OpenStudy (anonymous):

A couple wants to start a family. A doctor tells the couple based on many factors, the probability distribution of possible births is given. G is for girl and B is for boy. Assume the probabilities don't change for any reason. x G B P(X=x) 5/9 3/8 What is the probability that within the first 8 births, the family has at least 2 boys?

OpenStudy (anonymous):

remarkable that the probabilities do not add to one. is there some other choice besides boy or girl, or is there a typo?

OpenStudy (anonymous):

Um... first of all...

OpenStudy (anonymous):

the first 8?? jesus they need a tv

OpenStudy (anonymous):

P(transvestite children) = 5/72?

OpenStudy (anonymous):

this is the probability is: 1-p(0)-p(1) where \[p(X=x)=\left(\begin{matrix}8 \\ x\end{matrix}\right).p^{x}.(1-p)^{8-x}\]

OpenStudy (anonymous):

actually transvestite would count as male i think. maybe hermaphrodite

OpenStudy (anonymous):

I don't understand biology. =(

OpenStudy (anonymous):

question makes no sense, so i would not complete it

OpenStudy (anonymous):

maybe it is supposed to be \(\frac{5}{8}\) and \(\frac{3}{8}\) respectively

OpenStudy (anonymous):

I mean, maybe it's just a really gender-progressive question.

OpenStudy (anonymous):

she type it wrong 5/8

OpenStudy (anonymous):

It's for a stats class, and it's probability. it DOES make sense, I just don't know how to set it up. It's not typed wrong.

OpenStudy (anonymous):

well, maybe my teachers a retard. i guess that makes sense.

OpenStudy (anonymous):

no it does not. the probability has to add to one, because the complete sample space is G, B for each child

OpenStudy (anonymous):

Pretty box... the stats you gave mean that 5 out of 72 times, the child will be neither male nor female.

OpenStudy (anonymous):

we can pretend to complete the problem in any case. at least two boys means not one boy and not two boys probability of no boys in 8 births is \[(\frac{5}{8})^8\] probability of one boy is... well here is where we get stuck

OpenStudy (anonymous):

You mean not one boy and not 0 boys. But yes, I like this method.

OpenStudy (anonymous):

yeah that is what i mean. but the problem here is that \(1-p\) i don't know what to use

OpenStudy (anonymous):

Let's just pretend 5/8 G and 3/8 B.... Oi.

OpenStudy (anonymous):

probability of getting a boy is \(\frac{3}{8}\) so probability of not getting a boy should be \[1-\frac{3}{8}=\frac{5}{8}\] but it seems to have morphed in to \[\frac{5}{9}\] so it is not clear what number to use

OpenStudy (anonymous):

well then we can do it no boys = all girls probability is therefore \[(\frac{5}{8})^8\] one boy, seven girls probability is \[8\times (\frac{5}{8})^7\times \frac{3}{8}\] reasoning you have one boy, 7 girls, and 8 different ways to do this (first child is a girl, second child is a girl etc

OpenStudy (anonymous):

at least 2 boys means not all girls and not 7 girls so compute \[1-(\frac{5}{8})^8-8\times(\frac{5}{8})^7\times \frac{3}{8}\]

OpenStudy (anonymous):

thank you, that makes sense and i should be able to follow this pattern to figure out the other questions. <3